OpenStudy (skaaterchiick):

The solution to 4x + 7 - 9 = -14 is ??

imqwerty (imqwerty):

1 Simplify 4x+7−9 to 4x−2 4x−2=−14 2 Add 2 to both sides 4x=−14+2 3 Simplify −14+2 to −12 4x=−12 4 Divide both sides by 4 x=−12/4 5 Simplify 12/4 to 3 x=−3

OpenStudy (skaaterchiick):

Thanks

imqwerty (imqwerty):

np